# Cosmic Ray Magnetohydrodynamics: A New Two-Moment Framework with Numerical Implementation

### v1

@@ 3 Numerical Implementation
For notational convention, all the quantities without specifier are by default in lab frame hereafter.
3.1 Governing Equations
@@
$+\left[(\sigma_{+}^{L}-\sigma_{-}^{R})(u_{\parallel}+v_{A})+(\sigma_{-}^{L}-\sigma_{-}^{R})(u_{\parallel}-v_{A})\right]c\Delta P_{\rm cr}$
@@
Equation 50 and Equation 51 derived from the last section form our CR module, where $\mathcal{E}_{\rm cr}$, $\vec{F}_{\rm cr}$ and $\mathbf{P}_{\rm cr}$ are lab-frame CR energy density, energy flux and pressure tensor. $\sigma_{\pm}^{L,R}$ are scattering coefficients due to Alfvén waves, $\sigma_{\Omega}$ is a shorthand for $\Omega_{\rm cr}/c^{2}$, which is physically a large parameter, $\Delta P_{\rm cr}$ is the CR pressure anisotropy to be supplied by the user as a function of environmental parameters, $A$ is a prefactor that accounts for the average gamma factor $\left<\gamma\right$> in Equation 40 (for a 1GeV proton, $\gamma=1.06$ and thus $A\lesssim 1$), and $V_{m}$ is the reduced speed of light, which should not alter the simulation result as long as $V_{m}$ is significantly larger than any other signal speed (Skinner & Ostriker, 2013; Jiang & Oh, 2018). For most simulations presented in this paper, we adopt $V_{m}=100$ while other characteristic speeds remain of order unity. In all cases, we ensure that $V_{m}$ is much greater than any other characteristic speed.
3.2 HLLE Riemann Solver
@@
To update conserved variables in a cell $(k,j,i)$ by a finite-volume scheme, we first reconstruct the left $(\vec{q}_{L})$ and right $(\vec{q}_{R})$ states by the piecewise linear (second-order) reconstruction (Stone et al., 2008), and construct the HLLE fluxes $\mathbf{F}^{\rm HLLE}$ at cell interfaces. For example the HLLE flux at the (k,j,i1/2)(k,j,i-1/2) interface is:
@@
where $\mathbf{f}_{x}$ is the $x$ component (first column) of the flux tensor $\mathbf{f}$ calculated by substituting the matrix elements with corresponding left and right states, and $V^{+}$ and $V^{-}$ are the maximum and minimum wave speeds employed in the HLLE solver. Taking the transport terms split away from the source terms, the eigenvalues of hyperbolic Equation 52 give signal speeds $V^{\pm}=\pm V_{m}/\sqrt{3}$. However, as pointed out in Audit et al. (2002); Jiang et al. (2013), such large signal speed can lead to excessive numerical diffusion, particularly in the regime dominated by scattering opacity. Inspired by numerical studies for RHD, Jiang & Oh (2018) introduced a smooth transition between the free-streaming wave speed $V_{m}/\sqrt{3}$ and the diffusion-limit speed $V_{m}/\tau$ (where $\tau$ is the optical depth) to reduce the signal speed thereby numerical diffusion.
@@
where $\sigma$ stands for the summation of all the four scattering coefficients here, $\Delta x$ is the cell length in the direction being calculated, $C_{\rm max}\equiv\sqrt{C_{s}^{2}+C_{\rm cr}^{2}+v_{A}^{2}}$ with $C_{s}^{2}\equiv\gamma_{\rm gas}P_{\rm therm}/\rho$ and $C_{\rm cr}^{2}\equiv(4/3)P_{\rm cr}/\rho=\gamma_{\rm cr}P_{\rm cr}/\rho$ respectively being the gas sound speed and CR sound speed. The factor 1.5 in the last two cases serves as a safety factor. The detailed derivation explaining the transition in wave speeds for different regimes and the reasoning for the safety factor can be found in Appendix A.
3.3 Source Terms
@@
Although the implementation of explicit source terms is straightforward, we caution that it must be properly combined with the implicit integrators so that the overall algorithm can become more accurate than a first-order method (Huang & Bai, 2022). We denote the conserved CR variables after updating explicit source terms as $\mathcal{E}_{\rm cr}^{*}$, $\vec{F}_{\rm cr}^{*}$ and $\vec{q}^{*}$, which includes contributions from both advection and explicit source terms. We combine them with the implicit integrators described below.
As scattering is primarily along magnetic field lines for the implicit part, in each cell we first rotate the coordinate system by a rotation matrix $\mathbf{R}$ as in Jiang & Oh (2018):
@@
With these prerequisites, below we introduce our implicit schemes where the design of algorithm directly follows the dust-gas drag module in Huang & Bai (2022).
3.3.1 Backward Euler Method
@@
3.3.2 Implicit van Leer 2 (VL2) method
@@
$\vec{q}^{(n+1/2)}=\left(\mathbf{I}-\Delta t\mathbf{S}^{(n)}/2\right)^{-1}\left(\vec{q}^{(n)}+\Delta\vec{q}_{\rm af,src}^{(n)}\right)$ (63)
@@
Stage 2: use the updated quantities at the $n+1/2$ step to evolve the system from step $n$ to $n+1$. Here we directly list the eventual scheme while the detailed derivation can be found in the Appendix B [^5] The derivation in Huang & Bai (2022) was based on a general form of source terms, while our situation is a special case where the implicit source terms linearly depend on the conservative variables. of Huang & Bai (2022):
$\vec{q}^{(n+1)}-\vec{q}^{(n)}$ (64)
$=\mathbf{\Lambda}^{-1}\left(\mathbf{I}-\Delta t\mathbf{S}^{(n+1/2)}/2\right)\Delta t\left(\mathbf{S}^{(n+1/2)}\vec{q}^{(n)}+\Delta\vec{q}_{\rm af,src}^{(n+1/2)}\right)$
@@
$\mathbf{\Lambda}=\mathbf{I}-\left(\mathbf{I}-\Delta t\mathbf{S}^{(n+1/2)}/2\right)\Delta t\mathbf{S}^{(n+1/2)}$ (65)
@@
3.4 Stability Criterion for Time Step
It is well known that for equations describing a quantity streaming down its gradient, grid-scale oscillations can rapidly grow and overwhelm the true solution. We indeed observe this behavior in our 1D streaming tests (see more details in Section 4.1), where the scattering coefficients take the form given in Equation 68. As $\mathcal{E}_{\rm cr}$ streams down from a maixma, $\nabla P_{\rm cr}$ can flip sign due to an overshoot, which also flips the sign of $F_{\rm cr}$, then making the sign of $\nabla P_{\rm cr}$ change even more frequently. Consequently, numerical instability emerges.
To overcome this serious issue, Sharma et al. (2010a) regularized the one-moment CR transport equation by adding a numerical diffusion term. A disadvantage of this method is that it introduces artificial diffusion that can be hard to distinguish from the physical one, and also increases the computational expenses. Jiang & Oh (2018) instead removed this sign-flipping term from the RHS by combining it with the advection term on the LHS, and identified the modified source term as the effective scattering in the co-moving frame. This treatment is numerically convenient, but their formulation implicitly assumes that the scattering mechanism is solely subject to the CRSI and there is only one direction for Alfvén waves to propagate. Therefore, their method is inherently difficult to extend to the more general form we aim to consider in this work.
@@ 3.4 Stability Criterion for Time Step
$\Delta t<\frac{\sqrt{3}}{1+4\sigma_{0,\rm st}}\frac{\Delta x}{V_{m}}.$ (66)
@@
When $\sigma_{0,\rm st}\lesssim 1$, the $\Delta t$ given here is similar to the one limited by the transport term $(\Delta t<\sqrt{3}\Delta x/V_{m})$, so that a realistic time step can already reflect true physics [^6] For ion-neutral damping in real ISM, $\sigma_{0,\rm st}$ is typically $\lesssim 10^{-2}$ (see Section 2.2.3 in Armillotta et al. (2021) for estimation reference), thus does not encumber computational expense. For non-linear Landau damping that actively acts in ionized gas, the equivalent $\sigma_{0,\rm st}$ could be 3-4 orders of magnitude larger, thus limiting $\Delta t$. But in general it should still be affordable.. We caution that in principle this criterion is not aligned with other scattering physics (e.g., waves driven by CRSI but dissipated by non-linear Landau damping); however, we find that practically it also turns out to be applicable. We adopt this time step criterion in all simulations with streaming-dominated scattering coefficients (except for Figure 2 and Figure 12 where we test this stability condition).
4 Numerical Tests of the CR subsystem
We demonstrate the performance of our CR-MHD module in Athena++ via a series of numerical tests. Similar to Jiang & Oh (2018), we first test the CR subsystem in this section where the CR feedback to MHD gas is turned off. Additionally, except in Section 4.6 we neglect the $\mathcal{E}_{\rm cr}$ source term so that the CR energy is conserved and the subsystem is easier to analyze. In other words, for Section 4.1-Section 4.5, $\mathcal{E}_{\rm cr}$ is purely transported by the CR energy flux $\vec{F}_{\rm cr}$:
@@ 4 Numerical Tests of the CR subsystem
4.1 1D CR Streaming (linear damping)
@@
$\sigma_{\pm}^{L,R}=\frac{\sigma_{0,\rm st}}{2}\frac{|\nabla P_{\rm cr}|}{v_{A}P_{\rm cr}}\Theta(\mp\nabla P_{\rm cr}\cdot\vec{b}),$ (68)
where $\Theta$ is the Heaveside function. For a linear damping mechanism with rate $\nu_{\mathrm{damp}}$, the prefactor $\sigma_{0,\rm st}\sim\Omega_{\mathrm{cr}}(\rho_{\mathrm{cr}}/\rho_{\rm ion})/\nu_{\mathrm{damp}}$; this can take on a range of values depending on the CR energy density and environmental parameters, which we treat as a constant here.
@@
$F_{\rm cr}=-\frac{(1+4\sigma_{0,\rm st})}{3\sigma_{0,\rm st}}v_{A}\mathcal{E}_{\rm cr}{\rm sgn}\left(\frac{\partial\mathcal{E}_{\rm cr}}{\partial x}\right)+\frac{4}{3}u\mathcal{E}_{\rm cr}$ (69)
@@
in regions with $|\nabla\mathcal{E}_{\rm cr}|$ far from 0. We thus set a triangular profile for the initial CR energy density $\mathcal{E}_{\rm cr}(t=0,x)=2-|x|$ so that $|\nabla\mathcal{E}_{\rm cr}|$ is constant. We set $v_{A}=1$ with $\vec{b}=\hat{x}$, and the background MHD gas is static with all fluid variables fixed since the CR feedback to gas is turned off. Plugging the steady-state $F_{\rm cr}$ back to the conservative $\mathcal{E}_{\rm cr}$ equation, we obtain traveling-wave equation except at $\partial\mathcal{E}_{\rm cr}/\partial x$ discontinuities. Consequently, during the evolution of the system, $|\nabla\mathcal{E}_{\rm cr}|$ is either $1$ at the propagating fronts, or 0 in the central flat part originating from the initial discontinuity. For the default test here we use $\sigma_{0,\rm st}=1$, and set $\Delta t=0.3\Delta x/V_{m}$ according to our stability criterion (Equation 66).
@@
Jiang & Oh (2018) calculated the analytical solution of this problem, and by comparing it with our simulations in the left panel of Figure 1, we see our numerical results match very well with the analytical solutions except for the central regions in the $F_{\rm cr}$ panel. In the central plateau where $\nabla\mathcal{E}_{\rm cr}=0$, the time derivative $\partial F_{\rm cr}/\partial t$ can no longer be neglected and the theoretical expectation above is not valid. To test that numerical diffusion produced by the advection term does not overwhelm true physical solutions, we do the same streaming test in a moving gas background, where the background gas uniformly moves at a constant velocity $u=1$. This background movement should only introduce an additional $4u\mathcal{E}_{\rm cr}/3$ term in the steady $F_{\rm cr}$ (Equation 69), causing the center of $\mathcal{E}_{\rm cr}$ profile to drift at $4u/3$ on top of the previous solution. In the right panel of Figure 1 we show numerical results for the moving gas background, which agrees well with the analytic steady solution. We also measure the $L1$ errors in $\mathcal{E}_{\rm cr}$ at $t=0.3$ in our simulations with varying resolutions, and show in the bottom panel that the errors roughly decrease at a rate $\propto N_{x}^{-1.34}$ where $N_{x}$ is the cell number. Although we use a second-order scheme for this test, the scattering coefficients $\sigma_{\pm}^{L,R}$ depend on $\mathcal{E}_{\rm cr}$ but are treated as fixed parameters here. As a result, the method is not fully implicit, and the actual spatial convergence lies between first and second order.
@@
Figure 1: Numerical results of 1D streaming tests from Section 4.1. Top left: $\mathcal{E}_{\rm cr}$ and $F_{\rm cr}$ in a static background gas $(u=0)$. Colored lines denote snapshots at increasing simulation times, while dashed curves show analytical solutions, validating agreement between numerical and theoretical results. Top right: Identical setup with a uniformly advected background gas $(u=1)$, introducing additional $F_{\rm cr}$ from gas motion. The $\mathcal{E}_{\rm cr}$ profile drifts at $4u/3$, consistent with expectations. Bottom: Spatial convergence of $L1$ errors $(\sum_{i=1}^{N_{x}}|\mathcal{E}_{\rm cr,i}-\mathcal{E}_{\rm cr,theory}|/N_{x})$ at $t=0.3$. Errors roughly scale as $\propto N_{x}^{-1.34}$.
@@
Numerical results are stable in our $\sigma_{0,\rm st}=1$ tests, but we can increase the dimensionless prefactor $\sigma_{0,\rm st}$ by $100$ times so that the original $\Delta t$ violates the stability condition Equation 66. For this experiment, shown in the left column of Figure 2, we see both $\mathcal{E}_{\rm cr}$ and $F_{\rm cr}$ are choppy in the central plateau instead of connecting the left and right fronts smoothly. We then reduce $\Delta t$ by a factor of 0.1, as shown in the middle panels, which still can not fully stabilize the system. With a further reduction of $\Delta t$ by a factor 0.1, so that the stability condition is satisfied, the right column shows that stability is again recovered. This result lends support to our stability criterion as a useful time-step constraint for mitigating numerical instabilities in scenarios involving streaming-driven scattering terms.

### v2

@@ III Numerical Implementation
For notational convention, all quantities without an explicit specifier are, hereafter, by default in lab frame.
III.1 Governing Equations
@@
$+\left[(\sigma_{+}^{L}-\sigma_{+}^{R})(u_{\parallel}+v_{A})+(\sigma_{-}^{L}-\sigma_{-}^{R})(u_{\parallel}-v_{A})\right]c\Delta P_{\rm cr}$
@@
Equation 50 and Equation 51 derived from the last section form our CR module, where $\mathcal{E}_{\rm cr}$, $\vec{F}_{\rm cr}$ and $\mathbf{P}_{\rm cr}$ are lab-frame CR energy density, energy flux and pressure tensor. $\sigma_{\pm}^{L,R}$ are scattering coefficients due to Alfvén waves, $\sigma_{\Omega}$ is a shorthand for $\Omega_{\rm cr}/c^{2}$, which is physically a large parameter, $\Delta P_{\rm cr}$ is the CR pressure anisotropy to be supplied by the user as a function of environmental parameters, $A$ is a prefactor that accounts for the average gamma factor $\left<\gamma\right$> in Equation 40 (for a 1GeV proton, $\gamma=1.06$ and thus $A\lesssim 1$), and $V_{m}$ is the reduced speed of light, which should not alter the simulation result as long as $V_{m}$ is significantly larger than any other signal speed (Skinner and Ostriker, 2013; Jiang and Oh, 2018; Hopkins et al., 2022b). For most simulations presented in this paper, we adopt $V_{m}=100$ while other characteristic speeds remain of order unity. In all cases, we ensure that $V_{m}$ is much greater than any other characteristic speed.
III.2 HLLE Riemann Solver
@@
To update the conserved variables in a cell $(k,j,i)$ by a finite-volume scheme, we first reconstruct the left $(\vec{q}_{L})$ and right states $(\vec{q}_{R})$ using piecewise-linear (second-order) reconstruction (Stone et al., 2008). The same procedure is also applied to $\Delta P_{\rm cr}$ and magnetic fields, which enter the HLLE fluxes through $P_{{\rm cr},ij}$. Using these reconstructed quantities, we then compute the HLLE fluxes at the cell interfaces to obtain their divergence. For example, the HLLE flux at the (k,j,i1/2)(k,j,i-1/2) interface is:
@@
where $\mathbf{f}_{x}$ is the $x$ component (first column) of the flux tensor $\mathbf{f}$ calculated by substituting the matrix elements with corresponding left and right states, and $V^{+}$ and $V^{-}$ are the maximum and minimum wave speeds employed in the HLLE solver. Taking the transport terms split away from the source terms, the eigenvalues of hyperbolic Equation 52 give signal speeds $V^{\pm}=\pm V_{m}/\sqrt{3}$. However, as pointed out in Audit et al. (2002); Jiang et al. (2013), such large signal speed can lead to excessive numerical diffusion, particularly in the regime dominated by scattering opacity. Inspired by numerical studies for RHD, Jiang and Oh (2018) introduced a smooth transition between the free-streaming wave speed $V_{m}/\sqrt{3}$ and the diffusion-limit speed $V_{m}/\tau$ (where $\tau$ is the optical depth) to reduce the signal speed thereby numerical diffusion.
@@
where $\sigma$ stands for the summation of all the four scattering coefficients here, $\Delta x$ is the cell length in the direction being calculated, $C_{\rm max}\equiv\sqrt{C_{s}^{2}+C_{\rm cr}^{2}+v_{A}^{2}}$ with $C_{s}^{2}\equiv\gamma_{\rm gas}P_{\rm therm}/\rho$ and $C_{\rm cr}^{2}\equiv(4/3)P_{\rm cr}/\rho=\gamma_{\rm cr}P_{\rm cr}/\rho$ respectively being the gas sound speed and CR sound speed. The factor 2 in the last two cases serves as a safety factor. The detailed derivation explaining the transition in wave speeds for different regimes and the reasoning for the safety factor can be found in Appendix A.
III.3 Source Terms
@@
Although the implementation of explicit source terms is straightforward, we caution that it must be properly combined with the implicit integrators so that the overall algorithm can become more accurate than a first-order method (Huang and Bai, 2022). We denote the conserved CR variables after updating explicit source terms as $\mathcal{E}_{\rm cr}^{*}$, $\vec{F}_{\rm cr}^{*}$ and $\vec{q}^{*}$, which includes contributions from both advection and explicit source terms. We combine them with the implicit integrators described below.
As scattering is primarily along magnetic field lines for the implicit part, in each cell we first rotate the coordinate system by a rotation matrix $\mathbf{R}$ as in Jiang and Oh (2018):
@@
With these prerequisites, below we introduce our implicit schemes where the design of algorithm directly follows the dust-gas drag module in Huang and Bai (2022).
III.3.1 Backward Euler method
@@
III.3.2 Implicit van Leer 2 (VL2) method
@@
$\vec{q}^{(n+1/2)}=\left(\mathbf{I}-\frac{\Delta t}{2}\mathbf{S}^{(n)}\right)^{-1}\left(\vec{q}^{(n)}+\Delta\vec{q}_{\rm af,src}^{(n)}\right)$ (63)
@@
Stage 2: use the updated quantities at the $n+1/2$ step to evolve the system from step $n$ to $n+1$. Here we directly list the eventual scheme while the detailed derivation can be found in the Appendix B [^5] The derivation in Huang and Bai (2022) was based on a general form of source terms, while our situation is a special case where the implicit source terms linearly depend on the conservative variables. of Huang and Bai (2022):
$\vec{q}^{(n+1)}-\vec{q}^{(n)}=$ (64)
$\mathbf{\Lambda}^{-1}\left(\mathbf{I}-\frac{\Delta t}{2}\mathbf{S}^{(n+1/2)}\right)\Delta t\left(\mathbf{S}^{(n+1/2)}\vec{q}^{(n)}+\Delta\vec{q}_{\rm af,src}^{(n+1/2)}\right)$
@@
$\mathbf{\Lambda}=\mathbf{I}-\left(\mathbf{I}-\frac{\Delta t}{2}\mathbf{S}^{(n+1/2)}\right)\Delta t\mathbf{S}^{(n+1/2)}$ (65)
@@
III.4 Stability Criterion for Time Step
It is well known that for equations describing a quantity streaming down its gradient, grid-scale oscillations can rapidly grow and overwhelm the true solution. We indeed observe this behavior in our 1D streaming tests (see more details in Section IV.1), where the scattering coefficients take the form given in Equation 68. As $\mathcal{E}_{\rm cr}$ streams down from a maxima, $\nabla P_{\rm cr}$ can flip sign due to an overshoot, which also flips the sign of $F_{\rm cr}$, then making the sign of $\nabla P_{\rm cr}$ change even more frequently, generating numerical oscillations. The magnitude of such oscillations may be amplified or suppressed as simulation evolves, depending on the chosen time step.
To overcome this serious issue, Sharma et al. (2010a) regularized the one-moment CR transport equation by adding a numerical diffusion term. A disadvantage of this method is that it introduces artificial diffusion that can be hard to distinguish from the physical one, and also increases the computational expenses. Jiang and Oh (2018) instead removed this sign-flipping term from the RHS by combining it with the advection term on the LHS, and identified the modified source term as the effective scattering in the co-moving frame. This treatment is numerically convenient, but their formulation implicitly assumes that the scattering mechanism is solely subject to the CRSI and there is only one direction for Alfvén waves to propagate. Therefore, their method is inherently difficult to extend to the more general form we aim to consider in this work.
@@ III.4 Stability Criterion for Time Step
$\Delta t<\frac{\sqrt{3}}{1+4\widetilde{\sigma}_{0,\rm st}}\frac{\Delta x}{V_{m}}.$ (66)
@@
When $\widetilde{\sigma}_{0,\rm st}\lesssim 1$, the $\Delta t$ given here is comparable to the one limited by the transport term $(\Delta t<\sqrt{3}\Delta x/V_{m})$, and therefore does not impose a significant computational cost for resolving CR transport. This regime is typical when CR scattering is mediated by CRSI-driven waves damped by ion–neutral collisions in ISM, for which equating Equation 43 and Equation 68 yields $\widetilde{\sigma}_{0,\rm st}^{\rm IN}\sim\Omega_{\rm cr}(\rho_{\rm cr}/\rho_{\rm ion})/\nu_{\rm damp}\lesssim 10^{-2}$ (see e.g. Armillotta et al., 2022, for representative ISM parameters).
For non-linear Landau damping acting efficiently in ionized gas, the magnitude of the corresponding scattering coefficient can exceed $\sigma_{\rm IN}$ by 3-4 orders of magnitude (for reference, see equation 39 and 40 of Hix et al., 2025). Accordingly, the equivalent dimensionless parameter $\widetilde{\sigma}_{0,\rm st}^{\rm NLL}$ satisfies $\widetilde{\sigma}_{0,\rm st}^{\rm NLL}/\widetilde{\sigma}_{0,\rm st}^{\rm IN}=\sigma_{\rm NLL}/\sigma_{\rm IN}\sim 10^{3-4}$, leading to a more restrictive $\Delta t$. Nevertheless, the resulting $\Delta t$ remains computationally manageable in many practical cases.
We caution that this criterion was derived for scattering coefficients of the form given in Equation 68 (CRSI driving balanced by ion–neutral damping). Although one can formally compute an equivalent $\widetilde{\sigma}_{0,\rm st}$ for other damping mechanisms, such as the non-linear Landau damping discussed above, there is no guarantee that the resulting $\Delta t$ will rigorously ensure numerical stability. In practice, however, our test indicates that the criterion remains approximately applicable (see Section IV.2).
We adopt this time step criterion in all simulations with streaming-dominated scattering coefficients (except for Figure 2 and Figure 12 where the stability condition is explicitly tested).
IV Numerical Tests of the CR subsystem
We demonstrate the performance of our CR-MHD module in Athena++ via a series of numerical tests. Similar to Jiang and Oh (2018), we first test the CR subsystem in this section where the CR feedback to MHD gas is turned off. Additionally, except in Section IV.6 we neglect the $\mathcal{E}_{\rm cr}$ source term so that the CR energy is conserved and the subsystem is easier to analyze. In other words, for Section IV.1-Section IV.5, $\mathcal{E}_{\rm cr}$ is purely transported by the CR energy flux $\vec{F}_{\rm cr}$:
@@ IV Numerical Tests of the CR subsystem
IV.1 1D CR Streaming (linear damping)
@@
$\sigma_{\pm}^{L,R}=\frac{\widetilde{\sigma}_{0,\rm st}}{2}\frac{|\nabla P_{\rm cr}|}{v_{A}P_{\rm cr}}\Theta(\mp\nabla P_{\rm cr}\cdot\vec{b}),$ (68)
where $\Theta$ is the Heaviside function.
Here the dimensionless prefactor $\widetilde{\sigma}_{0,\rm st}$ account for the factor $\sim\Omega_{\rm cr}(\rho_{\rm cr}/\rho_{\rm ion})/\nu_{\rm damp}$ in Equation 43. In reality, it can take on a range of values depending on the CR density and environmental parameters. For simplicity, here we treat $\widetilde{\sigma}_{0,\rm st}$ as a constant, encoding the overall amplitude on top of the key scaling $|\nabla P_{\rm cr}|/P_{\rm cr}$ that reflects the characteristics of streaming.
@@
$F_{\rm cr}=-\frac{(1+4\widetilde{\sigma}_{0,\rm st})}{3\widetilde{\sigma}_{0,\rm st}}v_{A}\mathcal{E}_{\rm cr}{\rm sgn}\left(\frac{\partial\mathcal{E}_{\rm cr}}{\partial x}\right)+\frac{4}{3}u\mathcal{E}_{\rm cr}$ (69)
@@
in regions with $|\nabla\mathcal{E}_{\rm cr}|$ far from 0. We thus set a triangular profile for the initial CR energy density $\mathcal{E}_{\rm cr}(t=0,x)=2-|x|$ so that $|\nabla\mathcal{E}_{\rm cr}|$ is constant. We set $v_{A}=1$ with $\vec{b}=\hat{x}$, and the background MHD gas is static with all fluid variables fixed since the CR feedback to gas is turned off. Plugging the steady-state $F_{\rm cr}$ back to the conservative $\mathcal{E}_{\rm cr}$ equation, we obtain traveling-wave equation except at $\partial\mathcal{E}_{\rm cr}/\partial x$ discontinuities. Consequently, during the evolution of the system, $|\nabla\mathcal{E}_{\rm cr}|$ is either $1$ at the propagating fronts, or 0 in the central flat part originating from the initial discontinuity. For the default test here we use $\widetilde{\sigma}_{0,\rm st}=1$, and set $\Delta t=0.3\Delta x/V_{m}$ according to our stability criterion (Equation 66).
@@
Jiang and Oh (2018) calculated the analytical solution of this problem, and by comparing it with our simulations in the left panel of Figure 1, we see our numerical results match very well with the analytical solutions except for the central regions in the $F_{\rm cr}$ panel. In the central plateau where $\nabla\mathcal{E}_{\rm cr}=0$, the time derivative $\partial F_{\rm cr}/\partial t$ can no longer be neglected and the theoretical expectation above is not valid. To test that numerical diffusion produced by the advection term does not overwhelm true physical solutions, we do the same streaming test in a moving gas background, where the background gas uniformly moves at a constant velocity $u=1$. This background movement should only introduce an additional $4u\mathcal{E}_{\rm cr}/3$ term in the steady $F_{\rm cr}$ (Equation 69), causing the center of $\mathcal{E}_{\rm cr}$ profile to drift at $4u/3$ on top of the previous solution. In the right panel of Figure 1 we show numerical results for the moving gas background, which agrees well with the analytic steady solution. We also measure the $L1$ errors in $\mathcal{E}_{\rm cr}$ at $t=0.3$ in our simulations with varying resolutions, and show in the bottom panel that the errors roughly decrease at a rate $\propto N_{x}^{-1.34}$ where $N_{x}$ is the cell number. Although we use a second-order scheme for this test, the scattering coefficients $\sigma_{\pm}^{L,R}$ depend on $\mathcal{E}_{\rm cr}$ but are treated as fixed parameters here. As a result, the method is not fully implicit, and the actual spatial convergence lies between first and second order.
@@
Figure 1: Numerical results of 1D streaming tests from Section IV.1. Top left: $\mathcal{E}_{\rm cr}$ and $F_{\rm cr}$ in a static background gas $(u=0)$. Colored lines denote snapshots at increasing simulation times, while dashed curves show analytical solutions, validating agreement between numerical and theoretical results. Top right: Identical setup with a uniformly advected background gas $(u=1)$, introducing additional $F_{\rm cr}$ from gas motion. The $\mathcal{E}_{\rm cr}$ profile drifts at $4u/3$, consistent with expectations. Bottom: Spatial convergence of $L1$ errors $(\sum_{i=1}^{N_{x}}|\mathcal{E}_{\rm cr,i}-\mathcal{E}_{\rm cr,theory}|/N_{x})$ at $t=0.3$. Errors roughly scale as $\propto N_{x}^{-1.34}$.
@@
Numerical results remain stable in our tests with $\widetilde{\sigma}_{0,\rm st}=1$. However, as $\widetilde{\sigma}_{0,\rm st}$ increases in Equation 68, the magnitude of overshoot around $\nabla P_{\rm cr}$ extrema also grows, leading to stronger numerical oscillations. This behavior demands a tighter time step constraint to gradually suppress the oscillations rather than let them amplify into instabilities, consistent with the stability condition given by Equation 66. To further test this criterion, we increase the dimensionless prefactor $\widetilde{\sigma}_{0,\rm st}$ by a factor of 100, such that the original timestep $\Delta t$ now violates the stability condition. In this experiment (left column of Figure 2), we see both $\mathcal{E}_{\rm cr}$ and $F_{\rm cr}$ are choppy in the central plateau instead of connecting the left and right fronts smoothly. We then reduce $\Delta t$ by a factor of 0.1, as shown in the middle panels, which still can not fully stabilize the system. With a further reduction of $\Delta t$ by a factor 0.1, so that the stability condition is satisfied, the right column shows that stability is again recovered. This result lends support to our stability criterion as a useful time-step constraint for mitigating numerical instabilities in scenarios involving streaming-driven scattering terms.